Julia's age is 4 less than 5 times Jack's age. If 3 times Jack's age is added to Julia's age, the sum is equal to twice Julia's age. Find their ages.
I came up the the equation of 5x-4+3x. It doesn't work. Help!

Answer by nerdybill(7384)

Julia's age is 4 less than 5 times Jack's age. If 3 times Jack's age is added to Julia's age, the sum is equal to twice Julia's age. Find their ages.
.
Assign variables to things you don't know.
Let x = jack's age
5x-4 = julia's age
.
then from:
"If 3 times Jack's age is added to Julia's age, the sum is equal to twice Julia's age." we get our equation:
(5x-4)+3x = 2(5x-4)
5x-4+3x = 10x-8
8x-4 = 10x-8
4 = 2x
2 = x (Jack's age)
5x-4 = 5(2)-4 = 10-4 = 6 (Julia's age)